If the sum of the first n terms of an AP is 4n-n square,what is the first term. What is the first two terms,second term.
Answer by amarjeeth123(574)

sum of the first n terms of an AP is 4n-n square=n(4-n)=(n/2)(8-2n)=(n/2)[6+(n-1)(-2)]
sum of the first n terms of an AP=(n/2)[2a+(n-1)d]
Comparing both we get a=3 and d=-2
First term a1=3
Second term a2=a+d=3-2=1